OpenStudy (anonymous):

In an AC circuit, the current I amperes is given by I=20sin4t, where t is the time in milliseconds. Find the maximum value of the current and the first instant at which this maximum current occurs. Okay, so I've actually got the answer, but I'm not sure how I got it /shot These are my steps: I=20sin3t i find the maximum point first -1

ganeshie8 (ganeshie8):

Your work for finding max value is correct

ganeshie8 (ganeshie8):

I = 20 is the max value you have got, plugging that in the current equation gives you : \[\large 20 = 20\sin (4t)\]

ganeshie8 (ganeshie8):

solve \(\large t\)

OpenStudy (anonymous):

right, hang on

OpenStudy (anonymous):

20=20sin(4t) 1=sin(4t) sin^-1(1)=4t t=90/4 =45/2

ganeshie8 (ganeshie8):

you want to use radians i guess

ganeshie8 (ganeshie8):

sin^-1(1) = 4t pi/2 = 4t pi/8 = t

OpenStudy (anonymous):

oh uh good point yes no wonder

ganeshie8 (ganeshie8):

btw, that a bit cheating

OpenStudy (anonymous):

is that? why?

ganeshie8 (ganeshie8):

\[\large \sin^{-1}(1) = 4t\]

ganeshie8 (ganeshie8):

sin is periodic with a period of 2pi, so the correct way to solve it is \[\large \frac{\pi}{2} + 2n\pi = 4t\]

ganeshie8 (ganeshie8):

assuming t >= 0, the first time max value occurs when n = 0

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

how is simply subbing in the 20 'cheating' though?

ganeshie8 (ganeshie8):

subbing 20 is not cheating, saying sin^-1(1) = 4t means, pi/2 = 4t is bit lazy

OpenStudy (anonymous):

but but it's just solving an equation :o and sin^-1(1) is a special angle, isn't it? so getting π/2 without working would be natural

ganeshie8 (ganeshie8):

sin^-1(1) could be pi/2 or 5pi/2 or 9pi/2 or.. anythng so when you solve an equation, you should never say sin^-1(1) = 4t means pi/2 = 4t

ganeshie8 (ganeshie8):

pi/2 is just one solution, there are infinitely many other solutions. its like saying x^2 = 1 means x = 1 but x could be -1 also right ?

OpenStudy (anonymous):

oh yeahh there are other angles too thanks ^^ actually sorry, if you wouldn't mind explaining one thing again, i don't really get the π/2+2nπ=4t working up there is 2nπ the frequency?

ganeshie8 (ganeshie8):

nope

ganeshie8 (ganeshie8):

period of sin(x) is 2pi right ?

OpenStudy (anonymous):

yep

ganeshie8 (ganeshie8):

that means the value of sin(x) repeats after "2pi" time right ?

OpenStudy (anonymous):

and it repeats 4 times?

ganeshie8 (ganeshie8):

suppose if the value of sin(x) is 1, then after 2pi seconds, its value will be 1 too

OpenStudy (anonymous):

wait, okay go on..

ganeshie8 (ganeshie8):